# ProbRadarM3F: mmWave Radar-based Human Skeletal Pose Estimation with Probability Map Guided Multi-Format Feature Fusion

### v2

@@ Abstract
[…]
Index Terms:
@@
I Introduction
Smart medical care is paramount in hospitals and nursing homes in today’s society. To improve healthcare outcomes and ensure patient safety, there is an increasing need for human pose monitoring. The Internet of Things (IoT) has become instrumental in this regard, providing continuous and efficient monitoring capabilities [1, 2]. As shown in Fig. 1, IoT with human activity monitoring is essential to improve the convenience, safety, and interactivity of medical care. Traditionally, monitoring systems have relied on RGB cameras to estimate human pose and track human activity [3]. However, these methods pose significant privacy concerns and can be less effective in low-light or obscured environments.
Recent advancements focus on reducing privacy concerns by using non-visual signal-based sensing methods [4, 5]. These signals do not contain human facial information or visual photographs, thus addressing the need for human privacy in IoT application scenarios [6, 7]. Millimeter-wave (mmWave) radar has emerged as a prominent technology due to its high resolution and cost-effectiveness. MmWave radar can accurately estimate human skeletal poses without capturing identifiable visual information, making it ideal for preserving privacy. This paper focuses on mmWave radar-based human skeletal pose estimation, a critical component in smart sensor systems for medical care or smart home IoT applications.
@@ I Introduction
Figure 1: Consider an illustration of a medical IoT application: In this scenario, if a person receiving care at home experiences a fall, it is immediately detected, triggering an alarm at a remote care center and prompting the automatic dispatch of an ambulance. Our paper focuses on utilizing mmWave radar to estimate the human skeletal pose, a critical sensor component in this example of a medical IoT application. This approach preserves the user’s privacy while still providing the necessary monitoring information to the care center.
MmWave radar, a robust yet low-cost sensor, has been applied for environment perception in advanced driver-assistance systems (ADAS) [8, 9, 10, 11, 12, 13] and cooperative intelligent transportation systems (C-ITS) [14] in recent years. Considerable frontier research on mmWave radar has focused on achieving human activity sensing tasks. However, significant challenges remain.
@@
In the classical radar processing pipeline, various levels of data representation exist. For human skeletal pose estimation, point clouds and heatmaps are the most widely used forms of data representation. However, point cloud generation relies on hand-crafted parameters, and the process itself loses Doppler information and some positional data, leading to difficulties in capturing finer joint features. Heatmaps are often generated using the Fast Fourier Transform (FFT) method [15], which can extract range, Doppler, and angle of arrival (AoA) information from the raw radar signal to create different types of heatmaps. However, much of the information in the heatmap is considered redundant and not valued. Recently, many pose estimation works have been realized through a point cloud $+$ neural network pipeline or a heatmap $+$ neural network pipeline. Some studies have focused on improving the deep learning capability of networks and algorithms to address the mmWave radar-based precision problem. However, the importance of capturing the information present in the radar signal itself is often neglected, leading to the loss of significant effective information.
@@
To address the inefficiencies in traditional heatmap-based methods, a new method is proposed, constructing an efficient feature extraction framework. Specifically, the method incorporates two feature extraction branches. One branch adheres to conventional procedures, utilizing the FFT method to generate heatmaps for extracting range, Doppler, azimuth, and elevation information from mmWave radar signals. The other branch introduces a probability map generation and positional encoding method, designed to enhance the utilization of positional information in radar signals. A model is also proposed for fusing probability-guided positional features with traditional heatmap features. The contributions of this work are summarized below:
@@
    A mmWave radar-based multi-format feature fusion model, ProbRadarM3F, is presented for human skeletal pose estimation. The model is designed for raw mmWave radar signals, with features extracted from two branches: the FFT branch and the ProbPE branch. In the ProbPE branch, probability maps are generated, introducing a positional encoding method. To the best of our knowledge, this study represents the first attempt to extract additional probabilistic and positional features from mmWave radar signals to aid in pose estimation using a probability map-guided positional encoding approach. By fusing features in different formats, the keypoint prediction accuracy is significantly increased. ProbRadarM3F can serve as a new baseline for research and application in radar position information exploitation.
@@
    To validate the effectiveness of ProbRadarM3F by fair comparison with [16], our model was built upon the model used in [16] and experiments were conducted on the same dataset, HuPR [16], which provides purely raw radar signals. The results demonstrate that the design of a probability map-guided positional encoding strategy greatly improves the recognition performance of the model, leading to a significant enhancement in the precision of human skeletal pose prediction.
@@
The subsequent sections of this paper are organized as follows: Section II provides a comprehensive review of the research and methodology of mmWave radar for indoor sensing applications, particularly for human skeletal pose prediction. Section III details the implementation of the proposed ProbRadarM3F model. In Section IV, experiments conducted with ProbRadarM3F on the HuPR dataset are presented, and the results are evaluated. The advantages of its pose prediction capabilities are demonstrated. Finally, Section V summarizes the approach and findings and suggests potential directions for future research.
@@
Figure 2: Illustration of our proposed ProbRadarM3F model. This model consists of two main branches: the FFT branch and the ProbPE branch. Different colored arrows represent data streams from different radars. The initial input is the raw ADC data from two vertically placed mmWave radars. After processing the raw radar data into a radar data cube, the FFT branch processes the data to extract features from range-Doppler-azimuth heatmaps and range-Doppler-elevation heatmaps based on 4D-FFT. The ProbPE branch applies positional encoding to extract features from generated radar probability maps. In the following module, the multi-frame and multi-format features are fused, and a cross- and self-attention module is introduced to generate skeletal pose estimation.
II Related Works
II-A Millimeter-Wave Radar Indoor Sensing
@@
The field of millimeter-wave (mmWave) radar for indoor sensing has gained substantial advancements in recent years, driven by the growing demand for non-invasive, privacy-preserving technologies for human presence and activity detection. MmWave radar offers unique advantages for indoor sensing applications. The mmWave frequency can penetrate non-metallic objects while providing high-resolution imaging capabilities [17]. Its utility is increasingly evident in through-wall sensing [18], human detection and recognition [19], and vital sign monitoring [20, 21, 22].
@@
Initially, researchers used radar signal processing to extract waveform features, combining them with target distance, orientation, altitude, and motion information to perform simple action classification tasks [23]. Later, more advanced technologies such as deep learning were introduced for more accurate indoor sensing projects based on basic mmWave radar data. For instance, Wang et al. proposed a sequence-to-sequence (seq2seq) 3D temporal convolutional network with a self-attention method to estimate human hand pose [24]. Yi et al. utilized a multi-person detection model based on long short-term memory (LSTM) to determine human presence and localize their positions using a single 60GHz mmWave radio [25]. Wu et al. pioneered the work of segmenting human silhouettes from millimetre-wave RF signals by locating human reflections in radar frames and extracting features from surrounding signals with a human detection module and from aggregated frames with an attention-based mask generation module [26].
@@
Although these studies have been highly successful, they have focused more on changes in the structure of deep learning networks and neglected how to extract more valid features from the radar signal. Given the inherent limitations of mmWave radar data, this oversight may result in the loss of valuable information during processing. Therefore, in this paper, we pay special attention to the position information contained in the radar signals, which is long unnoticed but rich in information. The positional feature is extracted and combined with traditional feature to depict human body poses.
II-B MmWave based Human Skeletal Pose Estimation
@@
Since the application demands both privacy and precision, the advent of mmWave radar technology has markedly expanded the horizons of human pose estimation and recognition. A notable research direction in human pose estimation is the recognition of skeletal pose. The advantage of human skeletal pose estimation lies in its ability to provide a simplified but informative representation of human pose and movements, facilitating efficient and accurate recognition of human activity. Skeletal pose prediction derived from mmWave radar signals features critical spatial and temporal information about human activity without the privacy concerns associated with optical imaging.
@@
RF-Pose [27] is one of the pioneering works that first considered human skeleton reconstruction. Zhao et al. proposed a deep network from FWCM signals to estimate coarse parts of the human skeleton under the supervision of visual information. They subsequently improved the model to successfully predict 14 human joints, including the head, neck, shoulders, elbows, wrists, hips, knees, and feet. RF-Pose and its follow-up work utilized expensive customized hardware systems. Nowadays, more studies are based on low-cost mass-produced industrial millimeter-wave radar and have demonstrated good performance. Ding et al. presented a kinematic-constrained learning architecture that incorporates kinematic constraints with neural network learning for skeleton estimation based on range-Doppler heatmaps [28]. Kong et al. performed a convolution operation on the Range-Doppler Profile to detect the corresponding ranges and designed a two-stream deep learning architecture to extract body shape and motion features to predict skeleton joint coordinates and reconstruct body posture [29]. Cao et al. developed a model that incorporates part-level range-Doppler maps for individual body parts with local kinematic constraints and global constraints for reconstructing the human skeleton [30]. In this work, we also take advantage of human skeletal pose. Our extracted positional features based on probability maps are beneficial in predicting human key joints positions and predicting human posture.
III Proposed Methods
@@
In this section, detailed descriptions of ProbRadarM3F for human skeletal pose estimation, built upon [16], with additional fusion of multi-format features from mmWave radar data, are provided. The structure of the model is illustrated in Fig. 2. After processing the raw signal from mmWave radar, features are extracted from the radar data through the FFT branch and the ProbPE branch. Multi-frame features in different formats from the two branches are fused. Finally, the coordinates of human skeletal joints are predicted from the fused features via a cross-self-attention mechanism.
@@
III-A FFT Branch: Radar Data Processing and Range-Doppler-Angle Map Generating
@@
While the azimuth information of the horizontally placed radar is retained, the azimuth information of the vertically placed radar is considered as elevation angle information for the overall data. Eventually, the range-Doppler-azimuth map from the horizontal radar and the range-Doppler-elevation map from the vertical radar are obtained. To combine dynamic information across varying chirps within the range-Doppler-angle maps, the features from multiple chirps are merged by employing the M-Net [31]. The M-Net module utilizes a neural network to fuse different chirp information within a frame, instead of the traditional method, to solve for Doppler velocities, thereby outputting the merged features of the frame.
III-B ProbPE Branch: Probability Map Generating and Positional Encoding
As shown in Fig. 2, in the ProbPE branch, the radar data cube is initially subjected to a two-dimensional FFT (2D FFT). The FFT is executed in the fast-time dimension (the dimension of digitized chirp samples) to obtain range information. Due to the significant impact of velocity frequency effects in the slow-time dimension, where multiple frames correspond to the same range unit, a FFT is also conducted in this dimension to obtain the Doppler frequency.
To minimize irrelevant information that could increase computational load or interfere with feature extraction, a constant false-alarm rate (CFAR) method [32] is employed. This method is crucial in distinguishing between targets and interference noise based on intensity differences. A two-dimensional CFAR (2D-CFAR) is applied to range-Doppler maps to select the elements containing more valid information. In the 2D-CFAR detection process, targets are detected as accurately as possible to reduce false alarms. Guard cells and reference cells are introduced near each detection cell, as shown in the lower left corner of Fig. 3. Cell Averaging CFAR (CA-CFAR) is a common variant of CFAR, which uses the following formula to calculate the threshold for each cell under examination:
@@ III-B ProbPE Branch: Probability Map Generating and Positional Encoding
Subsequently, probability maps are derived from the filtered range bins, which serve as a foundational step in understanding the spatial distribution of potential targets. The concatenation set for the filtered range bins from the two radars is taken separately. The phase differences among various antennas yield distinct spikes, indicating frequency components from different directions. Angle information is obtained by performing FFT on target data from distinct virtual antennas. Angle FFT is executed to derive azimuth from the horizontal radar and elevation from the vertical radar. An averaging calculation in the Doppler dimension is performed because the focus at this point is on probability and position information. For the non-overlapping range bins filtered out from two radars, the Doppler dimensions are averaged directly to fill the corresponding range-angle vectors, facilitating the subsequent multiplication operations. Thus, range-azimuth and range-elevation vectors from the two radars are extracted, respectively. These vectors contain information on the radar’s intensity at specific ranges and angles, effectively characterizing the probability of a target’s presence at a given location. Consequently, radar probability maps are constructed, as illustrated in Fig. 3.
Figure 3: An illustration of probability map generation. It starts by filtering out the range bin through 2D CFAR on the range-Doppler heatmaps. Angle information is then extracted from the selected range, producing range-azimuth and range-elevation vectors. These vectors are normalized, transposed, and multiplied to create probability maps, indicating the likelihood of the target appearing at a specific position. The probability map serves as a guide for the positional encoding method.
@@
where $(\mathrm{pos}_{\theta},\mathrm{pos}_{\phi})$ represents the position being computed and $d$ represents the dimension of the positional encoding vector. This formula intricately encodes positional information by encoding each coordinate, $\mathrm{pos}_{\theta}$ and $\mathrm{pos}_{\phi}$, into a unique 32-dimensional vector through the application of sine and cosine functions. This approach captures the essence of spatial variations across the probability maps.
@@
III-C Multi-format Feature Fusion and Estimation Head
The features obtained from both branches are based on single-frame radar data. During human body movement, both previous and subsequent frames can be used as a reference for the current frame’s action information. To better utilize temporal information, the multi-frame information is jointly processed, considering the inherent temporal continuity of human motion. Data from multiple frames before and after the target frame are integrated, thus harnessing richer temporal features.
To integrate multi-format spatio-temporal features, multiple 3D convolutional layers are used to aggregate information. These 3D convolutional blocks help extract and consolidate features across both spatial and temporal dimensions. For each spatial scale, a 3D convolution block aggregates the residual temporal information, yielding three-layer encoded features. In addressing the challenge of integrating multi-scale features from distinct processing branches, the output dimensions of encoded features across layers are standardized. This uniformity facilitates the direct summation of positional and probability features with features from the other branch. Unlike the conventional concatenation method, this direct addition method preserves spatial coherency, particularly for positional features. The fused features are obtained as follows:

### v3

# Senior Member, IEEE

@@ Abstract
Millimeter wave (mmWave) radar is a non-intrusive, privacy-preserving, and cost-effective device, shown to be a viable alternative to RGB cameras for indoor human pose estimation. However, the challenge lies in fully leveraging the reflected radar signals for accurate pose estimation. To address this major challenge, this paper introduces a probability map guided multi-format feature fusion model, ProbRadarM3F. This is a radar feature extraction framework using a traditional FFT method in parallel with a probability map based positional encoding method. ProbRadarM3F fuses the traditional heatmap features and the positional features, then effectively achieves the estimation of $14$ keypoints of the human body. Experimental evaluation on the HuPR dataset proves the effectiveness of $69.9\%$ in average precision (AP). The emphasis of our study is on utilizing position information in radar signals for estimating human skeletal pose. This provides direction for investigating other potential non-redundant information from mmWave radar.
{IEEEkeywords}
@@
Figure 1: Illustration of a medical IoT application: if someone receiving care at home falls accidentally, it can be immediately detected, triggering an alarm at a remote care center and prompting the automatic dispatch of an ambulance. Our research focuses on utilizing mmWave radar to estimate the human skeletal pose, a critical sensor component in this example of a medical IoT application. This approach preserves the user’s privacy while still providing the necessary monitoring information to the care center.
1 INTRODUCTION
Smart medical care is paramount in hospitals and nursing homes in today’s society. To improve healthcare outcomes and ensure patient safety, there is an increasing need for human pose monitoring. The Internet of Things (IoT) has become instrumental in this regard, providing continuous and efficient monitoring capabilities [1, 2]. As shown in Fig. 1, IoT with human activity monitoring is essential to improve the convenience, safety, and interactivity of medical care. Traditionally, monitoring systems have relied on RGB cameras to estimate human pose and track human activity [3, 4]. However, RGB-based solutions can be unreliable in low-light or occluded settings. More critically, cameras raise substantial privacy concerns in sensitive indoor environments such as nursing homes, bedrooms, and hospital wards. People are reluctant to accept devices that capture visual imagery because of the risk of data leakage. Consequently, using indoor cameras to monitor the pose of people who are in need, like the elderly or patients, is often considered unacceptable.
Due to these mentioned reasons, recent advancements focus on reducing privacy concerns by using non-visual signal-based sensing methods [5, 6]. These signals contain no human facial information or visual photographs, thus protecting human privacy in IoT application scenarios [7, 8]. Millimeter-wave (mmWave) radar has emerged as a prominent technology due to its high resolution and cost-effectiveness. MmWave radar can accurately estimate human skeletal poses without capturing identifiable visual information, making it ideal for preserving privacy. This paper focuses on mmWave radar-based human skeletal pose estimation, a critical component in smart sensor systems for medical care and smart home IoT applications.
@@ 1 INTRODUCTION
MmWave radar, a robust yet low-cost sensor, has been applied for environment perception in advanced driver-assistance systems (ADAS) [9, 10, 11, 12, 13, 14] and cooperative intelligent transportation systems (C-ITS) [15] in recent years. Considerable frontier research on mmWave radar has focused on achieving human activity sensing tasks. However, significant challenges remain. In the classical radar processing pipeline, various levels of data representation exist. For human skeletal pose estimation, point clouds and heatmaps are the most widely used forms of data representation. However, point cloud generation relies on hand-crafted parameters, and the process itself loses Doppler information and some positional data, leading to difficulties in capturing finer joint features. Heatmaps are often generated using the Fast Fourier Transform (FFT) method [16], which can extract range, Doppler, and angle of arrival (AoA) information from the raw radar signal to create different types of heatmaps. However, much of the information in the heatmap is considered redundant and not valued. Recently, many pose estimation works have been realized through a point cloud $+$ neural network pipeline[17] or a heatmap $+$ neural network pipeline. Some studies have focused on improving the deep learning capability of networks and algorithms to address the mmWave radar-based precision problem. However, the importance of capturing the information present in the radar signal itself is often neglected, leading to the loss of significant effective information.
@@
To address the inefficiencies in traditional heatmap-based methods, a new method is proposed, constructing an efficient feature extraction framework. The method incorporates two feature extraction branches, generating multi-format features, i.e., distinct feature sets derived from the same raw data using different extraction strategies: the frequency features from heatmaps, and the positional features guided by probability maps. Specifically, one branch adheres to conventional procedures, utilizing the FFT method to generate heat maps to extract range, Doppler, azimuth, and elevation information from mmWave radar signals. The other branch introduces a probability map generation and positional encoding method, designed to enhance the utilization of positional information in radar signals. Our model fuses features in two formats (probability-guided positional features and traditional heat map features) and takes into account the influence and help of the frames before and after the target frame in the fusion. The contributions of this work include:
@@
    We propose ProbRadarM3F, a mmWave radar-based multi-format feature fusion model built on the model in [18], for the estimation of human skeletal pose. The model is designed to utilize raw mmWave radar signals, with different formats of features extracted from two branches: the FFT branch and the introduced ProbPE branch.
@@
    The ProPE branch proposed in this paper is a novel approach to generate probability map-guided positional encoding, as an additional format of radar feature. To the best of our knowledge, this study represents the first attempt to extract additional probabilistic and positional features from mmWave radar signals to aid in pose estimation. By fusing features in different formats from multiple frames, the keypoint prediction accuracy is significantly increased. Thus, ProbRadarM3F can serve as a new baseline for research and application in radar position information exploitation.
  • $\bullet$
@@
    To validate the effectiveness of ProbRadarM3F, experiments were conducted on the same dataset as HuPR [18], providing purely raw radar signals. The results demonstrate that the design of a probability map-guided positional encoding strategy greatly improves the recognition performance, leading to a significant enhancement in the precision of human skeletal pose prediction.
@@
This paper is organized as follows: Section 2 provides a comprehensive review of the research and methodology of mmWave radar for indoor sensing applications, particularly for human skeletal pose prediction. Section 3 details the implementation of the proposed ProbRadarM3F model. In Section 4, experiments conducted with ProbRadarM3F on the HuPR dataset are presented, and the results are evaluated. The advantages of its pose prediction capabilities are demonstrated. Finally, Section 5 summarizes the approach and findings and suggests potential directions for future research.
@@
Figure 2: Illustration of our proposed ProbRadarM3F model. This model consists of two main branches, namely the FFT branch and the ProbPE branch. Different colored arrows represent data streams from different radars. The initial input is the raw ADC data from two vertically placed mmWave radars. After processing the raw radar data into a radar data cube, the FFT branch processes the data to extract features from range-Doppler-azimuth heatmaps and range-Doppler-elevation heatmaps based on 4D-FFT. The ProbPE branch applies positional encoding to extract features from generated radar probability maps. In the following module, the multi-frame and multi-format features are fused, and a cross- and self-attention module [18] is introduced to generate estimated skeletal pose.
2 RELATED WORK
2.1 Millimeter-wave Radar Indoor Sensing
@@
The field of millimeter-wave (mmWave) radar for indoor sensing has gained substantial advancements in recent years, driven by the growing demand for non-invasive, privacy-preserving technologies for human presence and activity detection. MmWave radar offers unique advantages for indoor sensing applications. The mmWave frequency can penetrate non-metallic objects while providing high-resolution imaging capabilities [19]. Its utility is increasingly evident in through-wall sensing [20], human detection and recognition [21], and vital sign monitoring [22, 23, 24, 25].
@@
Initially, researchers used radar signal processing to extract waveform features, combining them with target distance, orientation, altitude, and motion information to perform simple action classification[26]. Later, advanced technologies, e.g., deep learning, were introduced for more accurate indoor sensing projects based on basic mmWave radar data. For instance, Wang et al. proposed a sequence-to-sequence (seq2seq) 3D temporal convolutional network with a self-attention method to estimate human hand pose [27]. Yi et al. utilized a multi-person detection model based on long short-term memory (LSTM) to determine human presence and localize their positions using a single 60GHz mmWave radio [28]. Ding et al. developed a novel hybrid neural network model that explores multi-domain fusion of radar information by utilizing three kinds of 2D domain spectra. They combine 1D and 2D convolutional neural network and a recurrent neural network to capture rich features through multiple domains. [29] Wu et al. pioneered the work of segmenting human silhouettes from millimeter-wave RF signals by locating human reflections in radar frames and extracting features from surrounding signals with a human detection module and from aggregated frames with an attention-based mask generation module [30].
@@
Although these studies have been highly successful, they have focused more on changes in the structure of deep learning networks and neglected how to extract more valid features from the radar signal. Given the inherent limitations of mmWave radar data, this oversight may result in the loss of valuable information during processing. Therefore, in this paper, we pay special attention to the position information contained in the radar signals, which is long unnoticed but rich in information. The positional feature is extracted and combined with traditional features to depict human body poses.
2.2 MmWave based Human Skeletal Pose Estimation
@@
Since the application demands both privacy and precision, the advent of mmWave radar technology has expanded the horizons of human pose estimation and recognition. A notable research direction in human pose estimation is the recognition of skeletal pose. The advantage of human skeletal pose estimation lies in its ability to provide a simplified but informative representation of human pose and movements, facilitating efficient and accurate recognition of human activity. Skeletal pose prediction derived from mmWave radar signals features critical spatial and temporal information about human activity without the privacy concerns associated with optical imaging.
@@
RF-Pose [31] is one of the pioneering works that first considered human skeleton reconstruction. Zhao et al. proposed a deep network from Frequency Modulated Continuous Wave(FMCW) radar signals to estimate coarse parts of the human skeleton under the supervision of visual information. They subsequently improved the model to successfully predict 14 human joints, including the head, neck, shoulders, elbows, wrists, hips, knees, and feet. RF-Pose and its follow-up work utilized expensive customized hardware systems. Nowadays, more studies are based on low-cost mass-produced industrial millimeter-wave radar and have demonstrated good performance. Ding et al. presented a kinematic-constrained learning architecture that incorporates kinematic constraints with neural network learning for skeleton estimation based on range-Doppler heatmaps [32]. Kong et al. performed a convolution operation on the Range-Doppler Profile to detect the corresponding ranges and designed a two-stream deep learning architecture to extract body shape and motion features to predict skeleton joint coordinates and reconstruct body posture [33]. Cao et al. developed a model that incorporates part-level range-Doppler maps for individual body parts with local kinematic constraints and global constraints for reconstructing the human skeleton [34]. Cao et al. proposed a two-stage framework for task-specific feature purification, which initially employs an adversarial auto-encoder to extract interpersonal-independent features, followed by a feature disentanglement module to eliminate components irrelevant to the pose [35].
While these studies have laid an important foundation in the estimation of skeletal pose using mmWave radar, they focus primarily on high-cost hardware systems, require substantial visual information supervision, or rely on complex processing of heat maps and point clouds. Without exploring the integration of probabilistic and positional features in a unified framework, existing methods may limit their applicability to various practical scenarios. In this work, our proposed extracting positional features based on probability maps are highly beneficial in accurately predicting human key joints positions and reconstructing human skeletal pose. This innovative approach enhances the robustness and precision of skeletal pose estimation.
3 PROPOSED METHODS
@@
In this section, detailed descriptions of ProbRadarM3F for human skeletal pose estimation, built upon [18], with additional fusion of multi-format features from mmWave radar data, are provided. The structure of the model is illustrated in Fig. 2. After processing the raw signal from mmWave radar, multi-format features are extracted from the radar data through the FFT and ProbPE branches. Multi-frame features in different formats from the two branches are fused. Finally, the coordinates of human skeletal joints are predicted from the fused features via a cross-self-attention mechanism.
@@
3.1 FFT Branch: Radar Data Processing and Range-Doppler-Angle Map Generating
@@
While the azimuth information of the horizontally placed radar is retained, the azimuth information of the vertically placed radar is considered as elevation angle information for the overall data. Eventually, the range-Doppler-azimuth map from the horizontal radar and the range-Doppler-elevation map from the vertical radar are obtained. Following the pioneering work [18], to combine dynamic information across varying chirps within the range-Doppler-angle maps, the features from multiple chirps are merged by employing the M-Net [36]. The M-Net module utilizes a neural network to fuse different chirp information within a frame, instead of the traditional method, to solve for Doppler velocities, thereby outputting the merged features of the frame.
3.2 ProbPE Branch: Probability Map Generating and Positional Encoding
As shown in Fig. 2, in the ProbPE branch, the radar data cube is initially subjected to a two-dimensional FFT (2D FFT). The FFT is executed in the fast-time dimension (the dimension of digitized chirp samples) to obtain range information. Due to the significant impact of velocity frequency effects in the slow-time dimension, where multiple frames correspond to the same range unit, an FFT is also conducted in this dimension to obtain the Doppler frequency.
To minimize irrelevant information that increases computation or interferes with feature extraction, a constant false-alarm rate (CFAR) method [37] is employed. This method is crucial in distinguishing between targets and interference noise based on intensity differences. A two-dimensional CFAR (2D-CFAR) is applied to range-Doppler maps to select the elements containing more valid information. In the 2D-CFAR detection process, targets are detected as accurately as possible to reduce false alarms. Guard cells and reference cells are introduced near each detection cell, as shown in the lower left corner of Fig. 3. Cell Averaging CFAR (CA-CFAR) is a common variant of CFAR, which uses the following formula to calculate the threshold for each cell under examination:
@@ 3.2 ProbPE Branch: Probability Map Generating and Positional Encoding
Subsequently, probability maps are derived from the filtered range bins, which serve as a foundational step in understanding the spatial distribution of potential targets. The concatenation set for the filtered range bins from the two radars is taken separately. The phase differences among various antennas yield distinct spikes, indicating frequency components from different directions. The angle information is obtained by performing FFT on target data from distinct virtual antennas. Angle FFT is executed to derive azimuth from the horizontal radar and elevation from the vertical radar. An averaging calculation in the Doppler dimension is performed because the focus at this point is on probability and position information. For the non-overlapping range bins filtered out from two radars, the Doppler dimensions are averaged directly to fill the corresponding range-angle vectors, facilitating the subsequent multiplication operations. Thus, range-azimuth and range-elevation vectors from the two radars are extracted, respectively. These vectors contain information on the radar’s intensity at specific ranges and angles, effectively characterizing the probability of a target’s presence at a given location. Consequently, radar probability maps are constructed, as illustrated in Fig. 3.
Figure 3: Illustration of the probability map generation, which is a more detailed representation of the lower left part of Fig.2. It starts by filtering out the range bin through 2D CFAR on the range-Doppler heatmaps. Angle information is then extracted from the selected range, producing range-azimuth and range-elevation vectors. These vectors are normalized, transposed, and multiplied to create probability maps, indicating the likelihood of the target appearing at a specific position. The probability map serves as a guide for the positional encoding method.
@@
where $(\mathrm{pos}_{\theta},\mathrm{pos}_{\phi})$ represents the computed position and $d$ represents the dimension of the positional encoding vector. This formula intricately encodes position information by encoding each coordinate, $\mathrm{pos}_{\theta}$ and $\mathrm{pos}_{\phi}$, into a unique 32-dimensional vector through sine and cosine functions. This approach captures the essence of spatial variations across the probability maps.
@@
3.3 Multi-format Feature Fusion and Estimation Head
The features obtained from both branches are based on single-frame radar data. In human body movement, both previous and subsequent frames can be used as references for the current frame action information. To better utilize temporal information, the multi-frame information is jointly processed, considering the inherent temporal continuity of human motion. Data from multiple frames before and after the target frame are integrated, thus harnessing richer temporal features.
In our model, multi-format refers to the extraction and fusion of features from different representations of the radar data. Specifically, features from two branches utilize two distinct formats: the FFT branch captures frequency domain features from traditional heatmaps, while the ProbPE branch generates probability maps that highlight the positional features. By combining these features of various formats, our model leverages the complementary strengths of both branches.
To integrate multi-format spatio-temporal features, multiple 3D convolutional layers are used to aggregate information. These 3D convolutional blocks help extract and consolidate features across both spatial and temporal dimensions. For each spatial scale, a 3D convolution block aggregates the residual temporal information, yielding three-layer encoded features. In addressing the challenge of integrating multi-scale features from distinct processing branches, the output dimensions of encoded features across layers are normalized. This uniformity facilitates the direct summation of positional and probability features with features from the other branch. Unlike the conventional concatenation method, this direct addition method preserves spatial coherency, particularly for positional features. The fused features are obtained as follows: